The balance in an account grew from $5,032 to $12,145 in 15 years, when interest was compounded quarterly. What was the interest rate? 


1
Expert's answer
2021-07-15T08:23:16-0400

Solution:

Given, "P=\$ \ 5032, A=\$\ 12145, t=15\ years=60\ quarters"

Let rate of interest per annum be "r\%" compounded annually.

Then, "A=P(1+\dfrac{r}{400})^{60}"

"\Rightarrow 12145=5032(1+\dfrac{r}{400})^{60}
\\ \Rightarrow 2.413553=(1+\dfrac{r}{400})^{60}
\\ \Rightarrow 1.014793=1+\dfrac{r}{400}
\\ \Rightarrow \dfrac{r}{400}=0.014793
\\ \Rightarrow r=5.9172"

Thus, the rate of interest is 5.9172% per annum compounded annually.